OpenStudy (anonymous):

Perform the following division (find the quotient). (7 y - 28) / (-7)

OpenStudy (anonymous):

7(y-4)/7(-1)=(y-4)/(-1)=-y+4

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

really what is wrong?

OpenStudy (anonymous):

yea it is

OpenStudy (anonymous):

You can check by multiplying. Only other way you can out it is: 4-y

OpenStudy (anonymous):

i had that too and it was wrong brosef. thanks for trying though. it doesnt matter anymore i already got the problem wrong :( it wont let me go back

OpenStudy (anonymous):

It's sad. You always can discuss it with instructor, if you are completely sure that you are right and you can prove it by multiplication.